OpenStudy (anonymous):

Mason walked 1/3 mile before lunch and 5/12 mile after lunch. How many miles did he walk in all? A.2/5mi B.3/4mi c.11/12mi D.1 1/2

OpenStudy (hockeychick23):

ok for this you need to find a least common denominator, do you know how to do that?

OpenStudy (anonymous):

A

OpenStudy (hockeychick23):

Thats incorrect

OpenStudy (mindblast3r):

the answer is B

OpenStudy (anonymous):

Take 1/3 + 5/12 which gets you 6/15, then simplify that getting you 2/5.

OpenStudy (hockeychick23):

@Dustin303 thats incorrect

OpenStudy (anonymous):

you need common denominators first otherwise you get the wrong answer

OpenStudy (hockeychick23):

1/3(4/4)= 4/12 Now they both have a common denominator of 12 and you can add them together 4/12+5/12= 9/12= 3/4

OpenStudy (mindblast3r):

lololololol @ dustin

OpenStudy (anonymous):

I did the same thing

OpenStudy (mindblast3r):